# Observing Double White Dwarfs with the Lunar GW Antenna

### v1

@@ Abstract
Aims. In this work, we present the capabilities of LGWA to detect and localize short-period double white dwarfs in terms of sky locations and distances. The analysis is performed using a realistic spatial distribution of sources, merger rates, and binary-mass distributions derived from current population synthesis models.
@@
Results. The analysis indicates that LGWA could detect approximately $\sim\mathcal{O}(30)$ monochromatic galactic sources and $\sim\mathcal{O}(10)$ extragalactic mergers, demonstrating the unique potential of decihertz gravitational-wave detectors to access and characterize extragalactic DWD populations. This will open new avenues for understanding Type Ia supernova progenitors and the physics of DWDs.
@@ 1 Introduction
The Lunar Gravitational-Wave Antenna (LGWA) is a proposed Moon-based gravitational-wave (GW) detector with an observation band spanning from about 1 mHz to 1 Hz (Ajith et al., 2025; Harms et al., 2021). The working principle relies on the excitation of the Moon’s vibrational modes, in particular the quadrupolar eigenmodes, induced by the apparent tidal forces associated with the transit of a GW. The amplitude of the modes can be measured by an array of 4 seismometers (van Heijningen et al., 2023), from which it is possible to reject the seismic background and reconstruct the original waveform Harms (2022). The possible observation band is determined by both the lunar response to GW excitation and the instrumental detection noise (Cozzumbo et al., 2023). The LGWA will fill the gap in the dHz band, bridging between the space-detector LISA and future terrestrial GW detectors like the proposed Einstein Telescope (ET) and Cosmic Explorer (CE) in a very physics-rich frequency window (Ajith et al., 2025; Sedda et al., 2021).
Other proposed lunar-based GW detectors are the Gravitational-wave Lunar Observatory for Cosmology (GLOC), now renamed Laser Interferometer Lunar Antenna (LILA), and the Laser Interferometer On the mooN (LION), which would all consist of triangle-shaped laser interferometers (Jani & Loeb, 2020; Jani et al., 2025; Amaro-Seoane et al., 2021). In the dHz band, other proposed detectors are the Deci-hertz Interferometer Gravitational wave Observatory (DECIGO) with its precursor mission B-DECIGO (Kawamura et al., 2006, 2011; Sato et al., 2017; Isoyama et al., 2018), the Taiji, Tianqin and TianGO detectors (Luo et al., 2021; Hu & Wu, 2017; Luo et al., 2016; Kuns et al., 2020), the Big Bang Observer (BBO) and the Advanced Laser Interferometer Antenna (ALIA) as follow on missions to LISA (Crowder & Cornish, 2005). All the aforementioned, proposed dHz detectors are long-baseline laser interferometers.
@@
The scenario involving two CO WD is called double-degenerate (DD), contrary to the single-degenerate scenario (SD) in which the supernova results from the He and H accretion of a CO WD from a non-degenerate companion. We note that both channels are probably present, especially taking into account the phenomenological variety of SN Ia events. Both scenarios benefit from indirect evidence, even if the DD scenario would better explain the lack of H and He emission in SN Ia spectra (Leonard, 2007). In conclusion, there is still no definitive evidence of the progenitor’s nature (Ruiter & Seitenzahl, 2025). The quest for the progenitor is particularly important given the role of SN Ia as standard candles in the measurement of the Hubble constant $H_{0}$ (Riess et al., 1995, 2022; Pascale et al., 2024), and the tension with CMB measurements (Ade et al., 2016; Aghanim et al., 2020; Di Valentino et al., 2021). In this paper, $H_{0}=70\,\text{km s}^{-1}\text{Mpc}^{-1}$ is assumed.
@@
Observing GWs from extragalactic DWDs would provide a means of independently calibrating cosmic distances. By identifying the host galaxy and combining its redshift with the luminosity distance inferred from the GW signal, it becomes possible to estimate the Hubble constant. These systems can act as standard sirens (Maselli et al., 2020; Schutz, 1986; Krolak & Schutz, 1987), offering a novel and complementary approach to addressing the Hubble tension, similar to the method employed with GW170817 (Schutz, 1986; Abbott et al., 2017). A multimessenger detection of a DWD that bursts into a SN Ia would directly prove the presence of the DD formation channel and calibrate the SN Ia as standard candles. The opposite (GW detection without EM counterpart or EM detection without GW event) does not necessarily rule out the possibility of a DD event, as there are models in which the deflagration happens a long time after the coalescence of the two CO WD (Shen et al., 2012). Furthermore, the detection of a GW signal without an electromagnetic counterpart does not preclude the identification of the host galaxy. If the DWD is well-localized by the GW signal, it can enable a ”dark siren” measurement of the Hubble constant (Finke et al., 2021; Leandro et al., 2022; Gair et al., 2023a, b).
@@
While lower-frequency instruments like LISA are expected to detect approximately 10,000 to 20,000 (Korol et al., 2017) or more (Korol et al., 2022) resolvable inspiral signals from DWD systems in the Milky Way, and millions more contributing to a confusion-limited GW foreground at low frequencies (Korol et al., 2022), LGWA will be uniquely capable of observing the final inspiral phase and merger of extragalactic DWDs. In fact, the merging frequency for these systems matches the peak sensitivity of LGWA (sec. 2). A simplified estimate of the detector’s observational capability can be obtained by comparing the GW amplitude from a nearby DWD source with the expected power spectral density (PSD) of the instrument, as detailed in Appendix A. A similar method was adopted by Pichardo Marcano et al. (2025) in their analysis of GLOC and LGWA. However, assessing the realistic potential of LGWA requires more than analyzing a single system with fixed parameter. To reliably estimate detection rates and parameter estimation capabilities, these results must be placed within the context of a realistic, population-wide model of DWDs. In particular, the binary-mass distribution is highly non-trivial (see Section 3), shaped by complex stellar evolution processes and diverse star-formation histories. Assuming a simplified, approximately uniform mass distribution leads to a significant overestimation of the number of detectable systems.
@@
In this work, an efficient and reliable method is provided to simulate the rare but loud population of short-period DWDs (sec. 3), taking into account the absolute abundance weighted by the supernovae type Ia (SNIa) rate, and the detailed distribution of the DWD set in the parameter space arising from the main astrophysical phenomena. The simulated population is then analyzed in order to assess the LGWA performances in terms of detection and parameter estimation, with particular care with regards to the localization capabilities (sec. 4). Our analysis, based on the assumption that observed SN are produced by DWD merger, suggest that LGWA will be able to detect from a few to ten extragalactic events during the 10 years mission lifetime, however the uncertainty on this estimation is heavily affected by the merging frequency of the systems, which depends on the complex physics of the matter-dominated merging process.
@@ 2.1 Merging frequency for DWD systems
The small radius of WDs (order of magnitude of $10^{6}$ m) means that the GW-radiating phase is very long, before other effects such as magnetic braking, Roche overflow and tidal disruption overcome. In the following, the maximum GW frequency reached by a DWD system as a function of the WD masses is estimated considering different masses $m_{1}$, $m_{2}$, eccentricity $\varepsilon=0$ and WD radius $R_{\text{WD}}$. The assumption $\varepsilon=0$ is justified by the fact that close WD binaries exhibit highly circularized orbits, a consequence of the CE phase experienced during the evolution. Hydrodynamic simulations (Lorén-Aguilar et al., 2005) show that the merging occurs when the distance between the WDs is approximately $d\approx 2-3\leavevmode\nobreak\ \leavevmode\nobreak\ R_{\text{WD}}$. This limit is heavily dependent on the physics of the merging process, and thus is not easily to model with simple considerations. In the following, we consider two scenarios: the Roche scenario and the Contact scenario. The estimates corresponding to the Roche and Contact approximations represent the two extreme scenarios that encompass the range of realistic detection situations, which will become accessible as waveform models improve.
@@
The MR relation can also be simulated by the population synthesis code SeBa (Portegies Zwart & Verbunt, 1996; Toonen et al., 2012) to better account for low-density regime effects  [^1] An accurate description of the SeBa usage is provided in sec. 3.2, for now it is used only to simulate the MR relation.. Compared to the more complete SeBa MR relation, we verify that the simple theoretical prediction (eq. 2) with $\mu_{e}=2$ holds well (better than 1%) for masses $\gtrsim 0.5M_{\odot}$; at lower masses the complete degeneration approximation is not satisfied, making eq. 2 less reliable. However low mass systems are not relevant, due to the excessively low merging frequency. In the following eq. 2 will thus be used to approximate the MR relation. Note that the beginning of the Roche overflow is set by the least massive WD, and the GW waveform is not sharply cut at this frequency.
@@
The matter of two WD behaves as a rigid solid; the disruption happens at the contact of the two spherical WDs. In this case, the merging distance is given by $d=R(m_{1})+R(m_{2})$, with $R(M)$ from eq. 2. This assumption brings the risk of overestimating the signal-to-noise-ratio (SNR), however it represents a much more realistic assumption: the frequency cutoff for the Roche scenario is to be intended as a strongly conservative estimation of the maximum frequency, up to which the waveform approximants available today are reliable. In reality we can expect the waveform to reach and pass the contact frequency, even if strongly influenced by mass effects. To account for the range between these two extreme scenarios, in this paper we consider both the scenarios for the extragalactic DWD population. For the galactic population only the Roche scenario is applied, since it is already largely sufficient to assure a detection.
@@
Figure 1: The plot shows the dependence of the merging frequency with respect to the component masses for the Roche scenario (upper semiplane) and contact scenario (lower semiplane). The reported population (red for $m_{1}+m_{2}>M_{\text{Ch}}$ and gray for $m_{1}+m_{2}<M_{\text{Ch}}$) is presented in sec. 3, and corresponds to the simulated Milky Way DWD population.
Figure 1 shows the approximate maximum frequency reached before the merging as a color scale with white contours for some values of $f_{\text{max}}$. The Roche and contact scenarios are represented by the upper and the lower half-planes, respectively. The simulated galactic population that will be discussed in sec. 3 is overlaid in the plot, distinguishing the super-Chandrasekhar fraction (red) from the sub-Chandrasekhar fraction (gray). Note that, even if in principle a super-Chandrasekhar DWD can exist with a mass combination of $m_{1}>0.7\,M_{\odot}$, $m_{2}<0.7\,M_{\odot}$, the population extends mainly in the region where both masses are $0.7\,M_{\odot}$. In particular, the most populated region in the high-mass ($M_{\text{tot}}>M_{\text{Ch}}$) area is the ”stripe” defined by $m_{1}\in[0.7,1.4]\,M_{\odot}$, $m_{2}\approx 0.8\,M_{\odot}$; for simplicity, in the following, this strip is called ”super-Ch branch”. The CO DWDs, namely those with component masses approximately $0.5<M_{CO}/M_{\odot}<1.2$, which are the principal candidates for a SN Ia progenitor, merge at $0.05$ - $0.1$ Hz, namely the most sensitive frequency band of LGWA. At these frequencies, neither LISA nor ET have sufficient sensitivity to detect the DWD merger, making LGWA a uniquely capable instrument for its observation (Ajith et al., 2025). Note that the systems above the super-Ch branch are extremely rare. It is customary to use a binary defined by $m_{1}=m_{2}=1.2M_{\odot}$ or even $1.4M_{\odot}$ to quickly test the capabilities of a detector when considering DWD mergers. However, these results show that such simplified estimates have limited value, as they do not accurately reflect the true underlying mass distribution.
@@ 2.2 The binary system GW emission
The system is modeled by two masses $m_{1}$ and $m_{2}$ separated by a distance $d$ following quasi-circular orbits on the $xy$ plane: for DWDs, a treatment of the eccentricity $\varepsilon$ is not necessary (see sec. 3). In the Keplerian limit, the slow inspiral of the two point-masses follows a series of stationary circular orbits with frequency given by eq. 3. The energy radiated from the system via GW emission is (Peters & Mathews, 1963):
@@
A waveform template is used for the estimation of the detector response, chosen from a set of approximants from LALsimulation (part of LALsuite, LIGO Scientific Collaboration et al. 2018), a software package currently used by the LIGO/Virgo/KAGRA collaboration for the analysis of GW data. We use the template IMRPhenomD (Husa et al., 2016; Khan et al., 2016; Pürrer et al., 2023). The model is already implemented inside the Fisher matrix code used for this study, GWFish [^3] https://gwfish.readthedocs.io/(see Section 4.1) and the evaluation is sufficiently fast to enable a large population study. In the LGWA frequency band, the modeling is also very accurate. The template depends on a set of parameters: the two masses $m_{1}$, $m_{2}$, the luminosity distance, the inclination angle $\theta$, namely the angle between the observer and the total angular momentum of the system, the polarization angle $\psi$, the initial phase of the waveform, the position in the sky in the usual (ra, dec) coordinates, the time of arrival at the geocenter at the nominal merger (without considering the real merger, namely the time at which the Roche overflow occurs), and a frequency cutoff used to simulate both a system that doesn’t merge during the observation period, and for which the waveform is truncated, or the occurrence of the Roche overflow in merging systems. A description of these parameters can be found in the documentation of GWFish.

### v2

@@ Abstract
Aims. In this paper, we present the capabilities of LGWA to detect and localize short-period double white dwarfs in terms of sky locations and distances. The analysis employs realistic spatial distributions and merger rates, as well as binary-mass distributions informed by population-synthesis models.
@@
Results. The analysis indicates that, over 10 years of observation, LGWA could detect approximately 30 monochromatic Galactic sources and 10 extragalactic mergers, demonstrating the unique potential of decihertz gravitational-wave detectors to access and characterize extragalactic DWD populations. This will open new avenues for understanding Type Ia supernova progenitors and the physics of DWDs.
@@ 1 Introduction
The Lunar Gravitational-Wave Antenna (LGWA) is a proposed Moon-based gravitational-wave (GW) detector with an observation band spanning from about 1 mHz to 1 Hz (Ajith et al., 2025; Harms et al., 2021). The working principle relies on the excitation of the Moon’s vibrational modes, in particular the quadrupolar eigenmodes, induced by the apparent tidal forces associated with the transit of a GW. The amplitude of the modes can be measured by an array of four seismometers (van Heijningen et al., 2023), from which it is possible to reject the seismic background and reconstruct the original waveform (Harms, 2022). The possible observation band is determined by both the lunar response to GW excitation and the instrumental detection noise (Cozzumbo et al., 2023). The LGWA can fill the gap in the dHz band, bridging between the space-detector LISA and future terrestrial GW detectors like the proposed Einstein Telescope (Punturo et al., 2010; Branchesi et al., 2023) and Cosmic Explorer (Evans et al., 2021, 2023; Gupta et al., 2024) in a very physics-rich frequency window (Ajith et al., 2025; Arca Sedda et al., 2021).
Other proposed lunar-based GW detectors are the Laser Interferometer Lunar Antenna (LILA), and the Laser Interferometer On the mooN (LION), which all consist of triangle-shaped laser interferometers (Jani & Loeb, 2020; Jani et al., 2025; Amaro-Seoane et al., 2021). In the dHz band, other proposed detectors are the Deci-hertz Interferometer Gravitational wave Observatory (DECIGO) with its precursor mission B-DECIGO (Kawamura et al., 2006, 2011; Sato et al., 2017; Isoyama et al., 2018), the Taiji, Tianqin and TianGO detectors (Luo et al., 2021; Hu & Wu, 2017; Luo et al., 2016; Kuns et al., 2020), the Big Bang Observer (BBO) and the Advanced Laser Interferometer Antenna (ALIA) as follow on missions to LISA (Crowder & Cornish, 2005). All the aforementioned, proposed dHz detectors are long-baseline laser interferometers.
@@
The scenario involving two CO WD is called double-degenerate (DD), contrary to the single-degenerate scenario (SD) in which the supernova results from the He and H accretion of a CO WD from a non-degenerate companion. We note that both channels are probably present, especially taking into account the phenomenological variety of SN Ia events. Both scenarios benefit from indirect evidence, even if the DD scenario would better explain the lack of H and He emission in SN Ia spectra (Leonard, 2007). In conclusion, there is still no definitive evidence of the progenitor’s nature (Ruiter & Seitenzahl, 2025; Maoz & Mannucci, 2012). The quest for the progenitor is particularly important given the role of SN Ia as standard candles in the measurement of the Hubble constant (Riess et al., 1995, 2022; Pascale et al., 2025), and the tension with CMB measurements (Ade et al., 2016; Aghanim et al., 2020; Di Valentino et al., 2021).
@@
Observing GWs from extragalactic DWDs would provide a means of independently calibrating cosmic distances. By identifying the host galaxy and combining its redshift with the luminosity distance inferred from the GW signal, it becomes possible to estimate the Hubble constant. These systems can act as standard sirens (Maselli et al., 2020; Schutz, 1986; Krolak & Schutz, 1987), offering a novel and complementary approach to addressing the Hubble tension, similar to the method employed with GW170817 (Schutz, 1986; Abbott et al., 2017). A multimessenger detection of a DWD that bursts into a SN Ia would directly prove the presence of the DD formation channel and calibrate the SN Ia as standard candles. The opposite (GW detection without EM counterpart or EM detection without GW event) does not necessarily rule out the possibility of a DD event, as there are models in which the deflagration happens a long time after the coalescence of the two CO WD (Shen et al., 2012). Furthermore, the detection of a GW signal without an electromagnetic counterpart does not preclude the identification of the host galaxy. If the DWD is well-localized by the GW signal, it can enable a ”dark siren” measurement of the Hubble constant (Finke et al., 2021; Leandro et al., 2022; Gair et al., 2023).
@@
While lower-frequency instruments like LISA are expected to detect approximately 10,000 to 20,000 (Korol et al., 2017) or more (Korol et al., 2022) resolvable inspiral signals from DWD systems in the Milky Way, and millions more contributing to a confusion-limited GW foreground at low frequencies (Korol et al., 2022), LGWA will be uniquely capable of observing the final inspiral phase and merger of extragalactic DWDs. In fact, the merging frequency for these systems matches the peak sensitivity of LGWA (sec. 2). A simplified estimate of the detector’s observational capability can be obtained by comparing the GW amplitude from a nearby DWD source with the expected power spectral density (PSD) of the instrument, as detailed in Appendix A. A similar method was adopted by Marcano et al. (2025) in their analysis of LILA (former GLOC) and LGWA. However, assessing the realistic potential of LGWA requires more than analyzing a single system with fixed parameter. To reliably estimate detection rates and parameter estimation capabilities, these results must be placed within the context of a realistic, population-wide model of DWDs. In particular, the binary-mass distribution is highly non-trivial (see Sect. 3), shaped by complex stellar evolution processes and diverse star-formation histories. Assuming a simplified, approximately uniform mass distribution leads to a significant overestimation of the number of detectable systems.
@@
In this work, an efficient and reliable method is provided to simulate the rare but loud population of short-period DWDs (Sect. 3), taking into account the absolute abundance weighted by the supernovae type Ia (SNIa) rate, and the detailed distribution of the DWD set in the parameter space arising from the main astrophysical phenomena. The simulated population is then analyzed in order to assess the LGWA performances in terms of detection and parameter estimation, with particular care with regards to the localization capabilities (sec. 5). Our analysis, based on the assumption that observed SN are produced by DWD merger, suggest that LGWA will be able to detect from a few to ten extragalactic events during the ten-year mission lifetime, however the uncertainty on this estimation is heavily affected by the merging frequency of the systems, which depends on the complex physics of the matter-dominated merging process.
Throughout this paper, $H_{0}=70\,\text{km s}^{-1}\text{Mpc}^{-1}$ is assumed.
@@ 2.1 Merging frequency for DWD systems
The small radius of WDs (order of magnitude of $10^{6}$ m) implies that the GW-radiating phase is very long, before other effects such as magnetic braking, Roche overflow and tidal disruption take over. In the following, the maximum GW frequency reached by a DWD system as a function of the WD masses is estimated considering different masses $m_{1}$, $m_{2}$, eccentricity $\varepsilon=0$ and WD radius $R_{\text{WD}}$. The assumption $\varepsilon=0$ is justified by the fact that close WD binaries exhibit highly circularized orbits, a consequence of the CE phase experienced during the evolution. Hydrodynamic simulations (Lorén-Aguilar et al., 2005) show that the merging occurs when the distance between the WDs is approximately two to three times the WD radius. This limit is heavily dependent on the physics of the merging process, and thus is not easily to model with simple considerations. In the following, we consider two scenarios: the “Roche scenario” and the “Contact scenario”. The estimates corresponding to the Roche and Contact approximations represent the two extreme scenarios that encompass the range of realistic detection situations, which will become accessible as waveform models improve.
@@
The MR relation can also be simulated by the population synthesis code SeBa (Portegies Zwart & Verbunt, 1996; Toonen et al., 2012) to better account for low-density regime effects  [^1] An accurate description of the SeBa usage is provided in Sect. 3.2, for now it is used only to simulate the MR relation.. Compared to the more complete SeBa MR relation, we verify that the simple theoretical prediction (eq. 2) with $\mu_{e}=2$ holds well (better than 1%) for masses $\gtrsim 0.5M_{\odot}$; at lower masses the complete degeneration approximation is not satisfied, making eq. 2 less reliable. However low mass systems are not relevant, due to the excessively low merging frequency. In the following eq. 2 will thus be used to approximate the MR relation. Note that the beginning of the Roche overflow is set by the least massive WD, and the GW waveform is not sharply cut at this frequency.
@@
The matter of two WD behaves as a rigid solid; the disruption happens at the contact of the two spherical WDs. In this case, the merging distance is given by $d=R(m_{1})+R(m_{2})$, with $R(M)$ from eq. 2. This assumption brings the risk of overestimating the signal-to-noise-ratio (S/N), however it represents a much more realistic assumption: the frequency cutoff for the Roche scenario is to be intended as a strongly conservative estimation of the maximum frequency, up to which the waveform approximants available today are reliable. In reality we can expect the waveform to reach and pass the contact frequency, even if strongly influenced by matter effects. To account for the range between these two extreme scenarios, in this paper we consider both the scenarios for the extragalactic DWD population. For the Galactic population only the Roche scenario is applied, since it is already largely sufficient to assure a detection.
@@
Figure 1: Dependence of the merging frequency with respect to the component masses for the Roche scenario (upper semiplane) and contact scenario (lower semiplane). The reported population (red for $m_{1}+m_{2}>M_{\text{Ch}}$ and gray for $m_{1}+m_{2}<M_{\text{Ch}}$) is presented in Sect. 3, and corresponds to the simulated Milky Way DWD population.
Figure 1 shows the approximate maximum frequency reached before the merging as a color scale with white contours for some values of $f_{\text{max}}$. The Roche and contact scenarios are represented by the upper and the lower half-planes, respectively. The simulated Galactic population that will be discussed in Sect. 3 is overlaid in the plot, distinguishing the super-Chandrasekhar fraction (red) from the sub-Chandrasekhar fraction (gray). Note that, even if in principle a super-Chandrasekhar DWD can exist with a mass combination of $m_{1}>0.7\,M_{\odot}$, $m_{2}<0.7\,M_{\odot}$, the population extends mainly in the region where both masses are $0.7\,M_{\odot}$. In particular, the most populated region in the high-mass ($M_{\text{tot}}>M_{\text{Ch}}$) area is the ”stripe” defined by $m_{1}\in[0.7,1.4]\,M_{\odot}$, $m_{2}\approx 0.8\,M_{\odot}$; for simplicity, in the following, this strip is called ”super-Ch branch”. The CO DWDs, namely those with component masses approximately $0.5<M_{CO}/M_{\odot}<1.2$, which are the principal candidates for a SN Ia progenitor, merge at $0.05$ - $0.1$ Hz, namely the most sensitive frequency band of LGWA. At these frequencies, neither LISA nor ET have sufficient sensitivity to detect the DWD merger, making LGWA a uniquely capable instrument for its observation (Ajith et al., 2025). Note that the systems above the super-Ch branch are extremely rare. It is customary to use a binary defined by $m_{1}=m_{2}=1.2M_{\odot}$ or even $1.4M_{\odot}$ to quickly test the capabilities of a detector when considering DWD mergers. However, these results show that such simplified estimates have limited value, as they do not accurately reflect the true underlying mass distribution.
@@ 2.2 The binary system GW emission
The system is modeled by two masses $m_{1}$ and $m_{2}$ separated by a distance $d$ following quasi-circular orbits on the $xy$ plane: for DWDs, a treatment of the eccentricity $\varepsilon$ is not necessary (see Sect. 3). In the Keplerian limit, the slow inspiral of the two point-masses follows a series of stationary circular orbits with frequency given by eq. 3. The energy radiated from the system via GW emission is (Peters & Mathews, 1963):
@@
A waveform template is used for the estimation of the detector response, chosen from a set of approximants from LALsimulation (part of LALsuite, LIGO Scientific Collaboration et al. 2018), a software package currently used by the LIGO/Virgo/KAGRA collaboration. We use the template IMRPhenomD (Husa et al., 2016; Khan et al., 2016; Pürrer et al., 2023). The model is very accurate in the LGWA frequency band and is already implemented in the Fisher-matrix code used for this study, GWFish [^3] https://gwfish.readthedocs.io/(see Sect. 4.1). Furthermore, it enables sufficiently fast analyses, even for large-population studies. In the LGWA frequency band, the modeling is also very accurate. The template depends on a set of parameters: the two masses $m_{1}$, $m_{2}$, the luminosity distance, the inclination angle $\theta$, namely the angle between the observer and the total angular momentum of the system, the polarization angle $\psi$, the initial phase of the waveform, the position in the sky in the usual (ra, dec) coordinates, the time of arrival at the geocenter at the nominal merger (without considering the real merger, namely the time at which the Roche overflow occurs), and a frequency cutoff used to simulate both a system that doesn’t merge during the observation period, and for which the waveform is truncated, or the occurrence of the Roche overflow in merging systems. A description of these parameters can be found in the documentation of GWFish.